# Segment Anything Meets Point Tracking

### v1

@@ 4.1 Datasets
In the following subsections, we present an overview of the datasets used in our study. Section 4.1.1 provides a brief introduction to the Video Object Segmentation task and outlines the specific datasets we utilize for this task. Similarly, Section 4.1.2 discusses the Video Instance Segmentation task and the dataset associated with it.
4.1.1 Video Object Segmentation
Video Object Segmentation (VOS) refers to the process of segmenting a specific object across an entire video sequence. Semi-supervised VOS (also known as one-shot VOS or semi-automatic VOS) is the primary setting for VOS on which we evaluate our method. In this setting, the ground truth object mask of the first frame is provided, and the task is to predict the masks for subsequent frames. Alternatively, the first frame label can be a bounding box instead of a segmentation mask, or a set of points as is the case for our method. We evaluate our method on four VOS datasets: DAVIS $2016$, DAVIS $2017$ [27], YouTube-VOS $2018$ [38], and MOSE $2023$ [8].
DAVIS $\mathbf{2016}$ [27].
DAVIS $2016$ is a single-object VOS benchmark, consisting of $20$ highly diverse video sequences, each of which possesses well-annotated segmentation masks.
DAVIS $\mathbf{2017}$ [27].
A multi-object extension of its $2016$ version, DAVIS $2017$ includes $60$ videos in the training set and $30$ videos in the validation set, comprising a total of $197$ different objects. The video scenarios within this dataset are small but diverse.
YouTube-VOS $\mathbf{2018}$ [38].
YouTube-VOS $2018$ is a large-scale dataset collected from YouTube, comprising $3471$ training videos encompassing $65$ categories and $474$ validation videos with an additional $26$ unseen categories. The diversity in categories and the inclusion of seen and unseen classes allow for a comprehensive evaluation of a given model’s generalization capability.
MOSE $\mathbf{2023}$ [8].
MOSE $2023$ is a recently introduced dataset designed for multiple object segmentation and tracking in complex scenes. This dataset is replete with challenges such as the transient visibility of objects, the presence of minute or less noticeable entities, extensive occlusions, and scenes with a high object density. By design, each video in this dataset must contain multiple objects so that occlusions must be present, and objects must show sufficient motion, as opposed to being stationary or showing little movement.
Metrics.
We report the standard evaluation metrics for video object segmentation [8, 27, 38], including region similarity $\mathcal{J}$, contour accuracy $\mathcal{F}$, and their average, $\mathcal{J\&F}$.
4.1.2 Video Instance Segmentation
Video Instance Segmentation (VIS) is a task that combines object detection, instance segmentation, and object tracking across video frames, which aims to identify and segment each object instance over the whole video sequence. This is a much less explored task compared to VOS but has been gaining interest. We evaluate our method on the dense-video task of the UVO v1.01.01.01.0 [34] dataset.
@@ 4.1.2 Video Instance Segmentation
The Unidentified Video Objects (UVO) dataset is designed to recognize and segment all objects regardless of the categories, even those unseen during training, thereby focusing on VIS in the open world. Each video in UVO features on average 12.3 object annotations, a considerable increase from previous datasets having only 2 or 3 objects per video on average. UVO sources its videos from the Kinetics-400 [14] dataset and contains three different splits: FrameSet, VideoSparseSet, and VideoDenseSet. The VideoDenseSet consists of 3-second clips annotated densely at 30fps and tracked over time. The primary goal of VideoDenseSet is to study video open-world segmentation. Objects identifiable under COCO categories carry their respective COCO labels, while ambiguous objects or those outside the COCO taxonomy are labeled as “other”. This meticulous and exhaustive annotation structure makes the VideoDenseSet ideal for research areas that require an understanding of videos in a dense and comprehensive manner, such as robotics, autonomous driving, and augmented-reality applications.
Metrics.
We evaluate our method using standard evaluation metrics in image instance segmentation, adapted for video instance segmentation [38]. These include Average Precision (AP) and Average Recall (AR) IoU-based metrics. Given that each instance in a video comprises a sequence of masks, unlike image instance segmentation, IoU computation is carried out not only in the spatial dimensions but also in the temporal dimension. This implies that the sum of intersections at every single frame is divided by the sum of unions at every single frame. These metrics are generally computed on a per-category basis and subsequently averaged across all categories. However, we work with the class-agnostic version of UVO.
4.2 Implementation Details
Training Data.
For our experiments, we use pre-trained checkpoints provided by the respective authors for both PIPS [11] and SAM. PIPS is trained exclusively on a synthetic dataset, FlyingThings++ [11], derived from the FlyingThings [24] optical flow dataset. This dataset includes multi-frame amodal trajectories with synthetic occlusions caused by moving objects. SAM, on the other hand, has been trained on the large-scale SA-1B dataset, the largest image segmentation dataset to date, with over 1 billion masks on 11M licensed and privacy-respecting images. It is noteworthy that neither of these datasets includes video segmentation data, and they do not overlap with any of our evaluation data. This effectively positions our model in a zero-shot video segmentation setting.
Model Variations.
Our experiments led to two optimal model hyperparameters, distinguished as SAM-PT (without reinitialization) and SAM-PT-reinit (with reinitialization). These configurations were derived from our ablation study in Sec. 4.3. However, we found that using iterative refinement negatively impacted both SAM-PT and SAM-PT-reinit on the MOSE dataset, and likewise hindered SAM-PT-reinit on the YouTube-VOS dataset. Consequently, iterative refinement was deactivated for these specific datasets. For DAVIS, we additionally report results for replacing SAM with HQ-SAM [15] and denote the model variants as HQ-SAM-PT and HQ-SAM-PT-reinit. The HQ-SAM variants use 3 iterative refinement iterations instead of 12 iterations.
VOS Evaluation.
VIS Evaluation.
For evaluating our method on the VIS task, we leverage SAM’s automatic mask generation capacity to generate up to 100 mask proposals for the initial frame. We then propagate these proposed masks throughout the entire video sequence using our method. We evaluate TAM [40], a concurrent method we compare against, in the same manner. Our mask proposal generation process is currently simplistic and does not create any proposals for subsequent video frames. Consequently, it cannot identify objects that emerge in later frames, placing it at a disadvantage compared to VIS methods that are capable of doing so. Despite this limitation, our approach provides a consistent platform for comparing zero-shot methods in terms of how effectively they propagate diverse mask proposals from the first frame.
4.3 Ablation Study
We conducted detailed ablation experiments on the DAVIS $2017$ validation subset to validate various components and designs of SAM-PT. We employed SAM’s ViT-H as the backbone, for all tests. Each aspect was examined sequentially, integrating the optimal settings obtained from prior experiments. To ensure statistical soundness, multiple iterations of each experiment were carried out (between 4 and 12 runs per setup), with findings represented as mean and standard deviation across these runs.
While these results provide insight, there may be a risk of overfitting due to our limited validation dataset. While we endeavored to maintain a consistent evaluation protocol, future research should aim for a larger validation set, possibly derived from the YouTube VOS $2018$ train dataset, to mitigate this concern.
Figure 5: Query Point Sampling Ablation Study. We report mean $\mathcal{J\&F}$ scores along with the standard deviation on the validation subset of DAVIS $2017$ for different number of positive points per mask and different point selection methods. Using $8$ positive points per mask leads to a $40$-point performance boost compared to using a single point. Given $8$ positive points, K-Medoids and Shi-Tomasi perform comparably well as point selection methods.
Query Point Sampling
Fig. 5 illustrates that the number of positive points and the choice of point selection methods significantly influence performance. Using 8 points per mask showed a remarkable 40-point performance enhancement compared to a single point. This substantiates the argument that a single positive point is inadequate for prompting SAM as it often results in the segmentation of partial objects only. Among the point selection methods, K-Medoids and Shi-Tomasi produced comparable results, with a slight preference towards K-Medoids owing to its marginally higher mean score and resilience to the number of positive points per mask.
@@ 4.3 Ablation Study
Table 2: Point Tracker and SAM Configuration Ablation Study Results. Using the best parameters from the ablation study in Fig. 5, we report the mean performance (with standard deviation) on the validation subset of DAVIS $2017$ to study the impact of (a) different point trackers, (b) the number of negative points per mask, (c) the use of iterative refinement and (d) patch similarity filtering. We find that the best configuration uses PIPS as the point tracker, $1$ negative point per mask, $12$ refinement iterations, and no patch similarity filtering. PT: point tracker. NP: negative points per mask. IRI: iterative refinement iterations. PS: point similarity filtering threshold.
SAM-PT Configuration DAVIS $\mathbf{2017}$ Validation [27]
[…]
(a) point tracker
RAFT [30] 00 00 $63.0\pm 0.6$ $60.7\pm 0.6$ $65.4\pm 0.5$
SuperGlue [28] 00 00 $21.7\pm 2.8$ $19.6\pm 2.1$ $23.8\pm 3.4$
SuperGlue [28] 00 $3$ $28.4\pm 3.1$ $24.7\pm 2.4$ $32.0\pm 3.8$
TapNet [9] 00 00 $60.9\pm 0.2$ $58.2\pm 0.3$ $63.5\pm 0.2$
PIPS [11] 00 00 $\mathbf{72.3}\pm 1.2$ $\mathbf{70.4}\pm 1.3$ $\mathbf{74.3}\pm 1.1$ $\mathbf{+9.3}$
(b) negative points per mask
PIPS 00 00 $72.3\pm 1.2$ $70.4\pm 1.3$ $74.3\pm 1.1$
PIPS $1$ 00 $\mathbf{74.1}\pm 0.7$ $\mathbf{72.1}\pm 0.6$ $\mathbf{76.1}\pm 0.7$ $\mathbf{+1.8}$
PIPS $8$ 00 $74.0\pm 0.8$ $71.9\pm 0.8$ $76.0\pm 0.9$
PIPS $16$ 00 $73.4\pm 0.6$ $71.4\pm 0.6$ $75.3\pm 0.6$
PIPS $72$ 00 $72.2\pm 0.4$ $70.3\pm 0.4$ $74.0\pm 0.4$
(c) iterative refinement iterations
PIPS $1$ 00 $74.1\pm 0.7$ $72.1\pm 0.6$ $76.1\pm 0.7$
PIPS $1$ $1$ $75.7\pm 0.7$ $73.4\pm 0.7$ $78.1\pm 0.6$
PIPS $1$ $3$ $76.0\pm 0.6$ $73.4\pm 0.7$ $78.6\pm 0.7$
PIPS $1$ $12$ $\mathbf{76.3}\pm 0.6$ $\mathbf{73.6}\pm 0.6$ $\mathbf{78.9}\pm 0.6$ $\mathbf{+2.2}$
(d) patch similarity filtering
PIPS $1$ $12$ $\mathbf{76.3}\pm 0.6$ $\mathbf{73.6}\pm 0.6$ $\mathbf{78.9}\pm 0.6$ none
PIPS $1$ $12$ $0.002$ $72.7\pm 2.0$ $70.2\pm 1.8$ $75.2\pm 2.1$
PIPS $1$ $12$ $0.01$ $70.7\pm 2.0$ $68.3\pm 1.8$ $73.2\pm 2.1$
Point Tracking.
Tab. 2a shows that PIPS [11] demonstrated superior performance over TapNet [9], SuperGlue [28], and RAFT [30]. TapNet’s limitations stem from its lack of effective time consistency and its training on 256x256 images, which hampered its performance with higher-resolution images. SuperGlue, while proficient in matching sparse features across rigid scenes, grapples with effectively matching points from the reference frame in dynamic scenes, particularly under object deformations. RAFT, being an optical flow model, faced difficulties handling occlusions. Although PIPS’s prior use in our experiments may have offered some hyperparameter advantages, its superior performance is primarily attributable to its more robust design that emphasizes trajectory modeling over eight subsequent frames. This approach fosters the generation of coherent point trajectories and enhances occlusion detection.
Negative Points.
Tab. 2b highlights that incorporating negative points had a favorable impact, particularly in reducing segmentation errors when points deviated from the target object. The addition of negative points empowered SAM to better handle the point trackers’ failure cases, leading to improved segmentation and a 1.8-point enhancement over the non-use of negative points. Note that throughout all experiments, we always used the mixed point sampling method for sampling negative points which amounts to using random sampling when there is only one negative point per mask.
Iterative Refinement.
The iterative refinement approach contributed to higher-quality masks and mitigated the impact of artifacts in SAM’s output. Tab. 2c displays that this yielded an improvement of 2.2 points over the non-refinement approach.
Patch Similarity.
Our initial findings in Tab. 2d suggest that using patch similarity to filter unreliable tracking points was overly restrictive in our context, leading to substantial deletion of points. Although it did not prove beneficial in our current setup, this aspect certainly warrants further exploration, particularly in scenarios involving point re-initialization.
Reinitialization.
Fig. 6 presents the performance of different reinitialization variants. In Tab. 5 and Tab. 6, we also show it brings 2.5 and 2.0 points improvements on MOSE and UVO benchmarks respectively. The re-initialization process enhanced robustness against points falling off objects. By reinitializing all points based on the current mask prediction, we account for errors in point tracker outputs by discarding incorrect points and starting fresh from the current mask prediction. However, this assumes that we trust the currently outputted mask, which may not always be the case and sometimes leads to failures.
Figure 6: Reinitialization Ablation Study. Mean $\mathcal{J\&F}$ scores (with std. dev.) on the validation subset of DAVIS $2017$ for different reinitialization variants and configurations. The best result is achieved by reinitialization variant $A$, 12 refinement iterations, and 72 negative points per mask. The reinitialization variants differ in how the timestep at which the points get reinitialized is chosen, see Appendix A for more details. Although using reinitialization improves the performance only marginally on the validation subset, this strategy demonstrates substantial improvement on the MOSE $2023$ (Tab. 5) and UVO (Tab. 6) datasets.
In summary, our best-performing SAM-PT model employs K-Medoids for point selection with 8 points per mask, PIPS for point tracking, a single negative point per mask, and employs 12 iterations for iterative refinement without patch similarity filtering. Meanwhile, using reinitialization achieved optimum performance with 12 refinement iterations and 72 negative points per mask.
4.4 Comparison with State-of-the-art Methods
All reported results were computed with official tools or official evaluation servers. Sec. 4.4.1 reports Video Object Segmentation results, including qualitative results on unseen web videos. Sec. 4.4.2 reports Video Instance Segmentation results.
4.4.1 Video Object Segmentation
Table 3: Quantitative results in semi-supervised VOS on the validation subset of DAVIS $2017$.
DAVIS $\mathbf{2017}$ Validation [27]
@@ 4.4.1 Video Object Segmentation
AGSS [20] $67.4$ $64.9$ $69.9$
AGAME [13] $70.0$ $67.2$ $72.7$
AFB-URR [18] $74.6$ $73.0$ $76.1$
STM [25] $81.8$ $79.2$ $84.3$
SWEM [21] $84.3$ $81.2$ $87.4$
RDE [17] $86.1$ $82.1$ $90.0$
SwinB-DeAOT-L [41] $86.2$ $83.1$ $89.2$
XMem [4] $87.7$ $84.0$ $91.4$
@@
Painter [35] $34.6$ $28.5$ $40.8$
DINO [2] $71.4$ $67.9$ $74.9$
SegGPT [36] 75.675.6{75.6}75.6 72.572.5{72.5}72.5 78.678.6{78.6}78.6
PerSAM-F [45] $71.9$ $69.0$ $74.8$
SAM-PT (ours) $76.3\pm 0.6$ $73.6\pm 0.6$ ${78.9}\pm 0.6$
SAM-PT-reinit (ours) ${76.6}\pm 0.7$ ${74.4}\pm 0.8$ ${78.9}\pm 0.6$
HQ-SAM-PT [15] (ours) $\mathbf{77.2}\pm 0.5$ $74.7\pm 0.5$ $\mathbf{79.8}\pm 0.4$
HQ-SAM-PT-reinit [15] (ours) $77.0\pm 0.7$ $\mathbf{74.8}\pm 0.8$ $79.2\pm 0.6$
Performance Overview.
Our proposed method outperforms others that have not been trained on any video object segmentation data on the DAVIS $2017$ dataset, as reflected in Tab. 3. A mean $\mathcal{J\&F}$ score of $76.6$ points exceeds the PerSAM-F by $4.7$ points and the SegGPT generalist model by a single point. The experiments were repeated 8 times for statistical robustness, and we report the mean and standard deviation of our method’s performance.
We also outperform PerSAM-F on the YouTube-VOS $2018$ and MOSE $2023$ datasets, achieving mean scores of $67.0$ and $41.0$ as shown in Tabs. 4 and 5. However, with different mask training data, our performance falls short when compared to SegGPT on the two datasets.
(a) Good cases for SAM-PT
(b) Good cases for SAM-PT-reinit
(c) Failure cases for SAM-PT
Figure 7: Visualization of SAM-PT on the DAVIS 2017 Validation [27] dataset. Given the first-frame masks, we sample 8 positive points and either 1 (SAM-PT) or 72 (SAM-PT-reinit) negative points per object to initialize SAM-PT and SAM-PT-reinit. Circles denote positive points, crosses denote negative points, and red symbols (circle and cross) denote that the point was predicted to be occluded.

### v2

@@ 4.1 Datasets
We evaluate our method on four VOS datasets: DAVIS $2016$, DAVIS $2017$ [30], YouTube-VOS $2018$ [44], and MOSE $2023$ [13]. DAVIS $2017$ is also used in our interactive point-based video segmentation study. We additionally devise a VOS dataset from BDD100K [50]. For VIS, We evaluate our method on the class-agnostic dense video instance segmentation task of the UVO v1.01.01.01.0 [40] dataset. UVO v1.01.01.01.0 is a VIS dataset aiming for open-world segmentation, where objects of any category, including those unseen in training, are identified and segmented.
4.2 Implementation Details
Training Data. For our experiments, we use pre-trained checkpoints provided by the respective authors for the point trackers (PIPS [32], CoTracker [19], etc.) and SAM. PIPS and CoTracker have been trained exclusively on synthetic data, PIPS on FlyingThings++ [16] and CoTracker on TAP-Vid-Kubric [14]. SAM has been trained on the large-scale SA-1B dataset, the largest image segmentation dataset to date. HQ-SAM is further trained on the HQ-Seg-44k [20]. Noteworthy, none of these datasets contain video segmentation data, nor do they intersect with any datasets we use for evaluation, situating our model within a zero-shot setting.
Interactive Point-Based Video Segmentation. To assess the interactive capabilities of SAM-PT, we simulate user refinement of video segmentation results through point additions and removals. We compare three methods: a non-tracking approach using SAM alone, an online method making one pass through the video with a target IoU quality, and an offline method employing multiple passes that progressively aim at higher IoU quality. These methods are detailed in the Supplementary Material, which includes pseudocode for each simulation variant.
4.3 Ablation Study
Our ablation experiments on the DAVIS $2017$ validation subset assessed different aspects of SAM-PT’s design. Despite the valuable insights, we acknowledge the dataset’s limited scope in representing diverse and complex segmentation challenges, such as occlusions and varying environmental conditions, may constrain the generalizability of our ablation’s findings. Future investigations could benefit from a more varied validation set, potentially sourced from the YouTube-VOS $2018$ training dataset, to enhance robustness.
Our findings, detailed in Tab. 2, underscore SAM-PT’s adaptability across leading long-term point trackers – PIPS [16], TAPIR [15], and notably CoTracker [19], which excelled due to its precise point tracking and reliable occlusion predictions. In contrast, PIPS++ [53] lagged despite being a more recent iteration of PIPS[16], due to the lack of occlusion prediction which is important for the effective use of point tracking for segmentation. TapNet [14] struggled due to less effective temporal consistency and high-resolution inputs. Traditional methods such as SuperGlue [33] and RAFT [35], which, although proficient in their respective domains, either struggle with the dynamic and deformable aspects of video scenes or cannot handle occlusion, highlighting the specialized efficacy of long-term trackers in the video segmentation landscape.
Table 2: We report the mean performance and standard deviation across eight runs on the validation subset of DAVIS $2017$ to study the impact of different point trackers.
Point Tracker DAVIS 2017 Validation [30]
@@ 4.3 Ablation Study
SuperGlue [33] $28.4$±$\pm 3.1$ $\pm 2.4$ $32.0$±$\pm 3.8$
TapNet [14] $60.9$±$\pm 0.2$ $58.2$±$\pm 0.3$ $63.5$±$\pm 0.2$
RAFT [35] $63.0$±$\pm 0.6$ $\pm 0.6$ $65.4$±$\pm 0.5$
PIPS++ [53] $73.2$±$\pm 0.5$ $69.9$±$\pm 0.5$ $76.6$±$\pm 0.5$
PIPS [16] $76.3$±$\pm 0.6$ $73.6$±$\pm 0.6$ $78.9$±$\pm 0.6$
TAPIR [15] $76.7$±$\pm 0.3$ $73.8$±$\pm 0.4$ $79.7$±$\pm 0.3$
CoTracker [19] $\mathbf{77.6}$±$\pm 0.7$ $\mathbf{74.8}$±$\pm 0.7$ $\mathbf{80.4}$±$\pm 0.7$
In Tab. 3, we tested SAM-PT with various settings using PIPS as the tracker. We found that using eight positive points per object instead of just one improved our scores significantly by $33.4$ points because one point often wasn’t enough for unambiguously prompting SAM. Selecting points with K-Medoids was slightly better than random and matched Shi-Tomasi, giving a boost of $1.8$ points. Incorporating negative points besides positive points helped when trackers made mistakes, such as losing track of an object, by improving scores by another $1.8$ points. Adding iterative refinement smoothed out mask quality and fixed some errors, adding another $2.2$ points to our performance. We tried filtering out unreliable points with patch similarity, but this did not work well as it ended up removing too many points. Finally, although reinitializing points did not help significantly in the initial tests, it did show benefits on other datasets such as MOSE and UVO, helping us recover from tracker errors by discarding incorrect and adding fresh points as well as detecting that the object has disappeared and the tracking should be halted.
@@
Table 3: Ablation study on the validation subset of DAVIS $2017$ on the impact of different SAM-PT configurations using PIPS [16] as the point tracker. PSM: point selection method. PP: positive points per mask. NP: negative points per mask. IRI: iterative refinement iterations. PS: point similarity filtering. RV: reinitalization variant.
SAM-PT Configuration (using PIPS) DAVIS [30]
PSM PP NP IRI PS RV $\mathcal{J\&F}$ Gain
Random $1$ 00 00 $37.1$±$\pm 21.7$
Random $8$ 00 00 $70.5$±$\pm 1.4$ $+33.4$
K-Medoids $8$ 00 00 $72.3$±$\pm 1.2$ $+1.8$
Shi-Tomasi $8$ 00 00 $72.0$±$\pm 0.3$
Mixed $8$ 00 00 $70.6$±$\pm 0.8$
K-Medoids $8$ $1$ 00 $74.1$±$\pm 0.7$ $+1.8$
K-Medoids $8$ $1$ $12$ $76.3$±$\pm 0.6$ $+2.2$
K-Medoids $8$ $1$ $12$ $72.7$±$\pm 2.0$ none
K-Medoids $8$ $72$ $12$ A $76.8$±$\pm 0.7$ $+0.5$
K-Medoids $8$ $1$ $12$ B $76.1$±$\pm 0.4$
K-Medoids $8$ $1$ 00 C $75.5$±$\pm 0.7$
K-Medoids $8$ $1$ $12$ D $76.4$±$\pm 0.3$
Extended ablation experiments and discussions can be found in the Supplementary Material, including complete ablation results for PIPS and CoTracker, the choice of the SAM backbone, and SAM’s lightweight variants that reveal trade-offs between performance and inference speed.
4.4 Video Object Segmentation
Performance Overview. Our SAM-PT method, utilizing HQ-SAM and CoTracker, sets a new standard in zero-shot video object segmentation on the DAVIS $2017$ dataset with a mean $\mathcal{J\&F}$ score of $79.4$, outperforming SegGPT’s $75.6$, DINO’s $71.4$, and Painter’s $34.6$ as shown in Tab. 4. On the easier DAVIS $2016$ validation set, our method achieves $84.3$, surpassing SegGPT’s $82.3$, showcasing the strength of our approach even in less complex scenarios, as detailed in the Supplementary Material.
Table 4: Quantitative results on the DAVIS $2017$ [30] validation set for semi-supervised VOS. Performance is reported for different methods, including our SAM-PT and HQ-SAM-PT, with and without the reinitialization strategy (Reinit) and using different point trackers. Our method outperforms other zero-shot methods.
Method Tracker Reinit DAVIS $\mathbf{2017}$ Validation [30]
@@ 4.4 Video Object Segmentation
MiVOS [8] - - $84.5$ $81.7$ $87.4$
DeAOT [48] - - $86.2$ $83.1$ $89.2$
DEVA [10] - - $87.6$ $84.2$ $91.0$
XMem [7] - - $87.7$ $84.0$ $91.4$
@@
Painter [41] - - $34.6$ $28.5$ $40.8$
DINO [4] - - $71.4$ $67.9$ $74.9$
SegGPT [42] - - 75.675.6{75.6}75.6 72.572.5{72.5}72.5 78.678.6{78.6}78.6
SAM-PT (ours) PIPS [16] $76.3$±$\pm 0.6$ $73.6$±$\pm 0.6$ 78.978.9{78.9}78.9±$\pm 0.6$
PIPS [16] 76.676.6{76.6}76.6±$\pm 0.7$ 74.474.4{74.4}74.4±$\pm 0.8$ 78.978.9{78.9}78.9±$\pm 0.6$
CoTracker [19] 77.677.6{77.6}77.6±$\pm 0.7$ 74.874.8{74.8}74.8±$\pm 0.7$ 80.480.4{80.4}80.4±$\pm 0.7$
CoTracker [19] $77.4$±$\pm 1.0$ $74.5$±$\pm 1.0$ $80.3$±$\pm 1.1$
HQ-SAM-PT (ours) PIPS [16] $77.2$±$\pm 0.5$ $74.7$±$\pm 0.5$ $79.8$±$\pm 0.4$
PIPS [16] $77.0$±$\pm 0.7$ $74.8$±$\pm 0.8$ $79.2$±$\pm 0.6$
CoTracker [19] 79.4±$\pm 0.6$ 76.5±$\pm 0.6$ 82.3±$\pm 0.5$
CoTracker [19] $77.7$±$\pm 0.8$ $74.6$±$\pm 0.9$ $80.8$±$\pm 0.7$
For the YouTube-VOS $2018$ validation set, we achieve the highest performance among zero-shot methods with $76.2$ against SegGPT’s $74.7$ and Painter’s $24.1$, indicating robust generalizability across various video segmentation benchmarks (Tab. 5).
Table 5: Quantitative results in semi-supervised VOS on the validation subset of YouTube-VOS $2018$. Metrics are reported separately for “seen” and “unseen” classes, with $\mathcal{G}$ being the overall average score over the metrics.
Method Tracker Reinit YouTube-VOS $\mathbf{2018}$ Validation [44]
$\mathcal{G}$ $\mathcal{J}_{s}$ $\mathcal{F}_{s}$ $\mathcal{J}_{u}$ $\mathcal{F}_{u}$
(a) trained on video segmentation data
XMem [7] - - $86.1$ $85.1$ $89.8$ $80.3$ $89.2$
DeAOT [48] - - $86.2$ $85.6$ $90.6$ $80.0$ $88.4$
(b) not trained on video segmentation data (zero-shot)
Painter [41] - - $24.1$ $27.6$ $35.8$ $14.3$ $18.7$
SegGPT [42] - - 74.774.7{74.7}74.7 75.175.1{75.1}75.1 $\mathbf{80.2}$ 67.467.4{67.4}67.4 75.975.9{75.9}75.9
SAM-PT (ours) PIPS [16] $67.0{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.3}}$ $68.6{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$ $71.2{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.1}}$ $61.0{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.5}}$ $67.4{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.4}}$
PIPS [16] ${67.5}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$ ${69.0}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.4}}$ ${69.9}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.3}}$ ${63.2}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.4}}$ ${67.8}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.5}}$
CoTracker [19] $74.0{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.3}}$ $73.3{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$ $76.0{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$ $70.0{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.4}}$ $76.7{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.4}}$
CoTracker [19] $71.5{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.4}}$ $71.0{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.3}}$ $72.8{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.3}}$ $68.3{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.7}}$ $73.9{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.7}}$
HQ-SAM-PT (ours) CoTracker [19] $\mathbf{76.2}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.1}}$ $\mathbf{75.3}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.1}}$ $78.4{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$ $\mathbf{72.1}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$ $\mathbf{79.0}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$
In the semi-supervised VOS on BDD100K’s validation set, our method outperforms SegGPT for non-transient objects but also surpasses the fully-supervised XMem across nearly all object visibility durations. The detailed breakdown is provided in Tab. 6.
On the MOSE $2023$ validation set, our performance remains competitive with SegGPT, with exact figures available in the Supplementary Material.
(a) Successful segmentation cases for SAM-PT using 8 positive and 1 negative point.
(b) Successful segmentation cases for SAM-PT with 8 positive and 72 negative points and reinitialization enabled.
(c) Failure cases for SAM-PT where challenges such as occlusions and thin object structures lead to tracking errors.
Figure 5: Visualization of SAM-PT on DAVIS 2017 [30]. The method shows its capability to segment and track objects using the initial masks from the first frame, with circles denoting positive points and crosses negative points. Red symbols indicate occlusion prediction.